Assessment Scale for Written Work

E.Fleischmann, R.Huber, E. Pölzleitner, R. Vaupetitsch, K. Weißenböck

Adapted from: Jacobs, Holly L. et al., Testing ESL Composition: A Practical Approach, Newbury House, Rowley 1981

Please use according to the instructions on the reverse side

Content 20-18 Excellent to very good: Excellent to very good knowledge of subject; considerable variety of ideas or arguments; independent and thorough treatment of topic; relevant to assigned topic; accurate detail.

17-14 Good to Average: Adequate knowledge of subject; variety of ideas/arguments

above average; adequate treatment of topic; some detail, mostly relevant to topic. 13-10 Fair to Poor: Knowledge of subject is hardly adequate; limited variety of

ideas/arguments; sufficient treatment of topic; lacking detail or extraneous material included.

9-6 Very Poor: Knowledge of subject too little; hardly any or no variety of

ideas/arguments; insufficient treatment of topic, or not enough to evaluate.

Organization 20-18 Excellent to Very Good: Fluent expression; ideas clearly stated and supported; well organized; logically sequenced (coherent); connectives appropriately used (cohesive).

17-14 Good to Average: Somewhat choppy but main ideas stand out; logically sequenced;

connectives sometimes used inappropriately. 13-10 Fair to Poor: Choppy; loosely organized; main ideas are not always clear; connec-

tives often used inappropriately. 9-6 Very Poor: Non-fluent; ideas often confused or disconnected; lacks logical sequencing

or development.

Vocabulary 25-23 Excellent to Very Good: Wide range of vocabulary; accurate word/idiom choice

and usage; appropriate register. 22-18 Good to Average: Adequate range of vocabulary; occasional errors of word/idiom

form; choice and usage but meaning not obscured; register not always appropriate. 17-13 Fair to Poor: Limited range of vocabulary; frequent errors of word/idiom form,

choice and usage; meaning confused or obscured; inappropriate register. 12-8 Very Poor: Range of vocabulary too limited; essentially translation; meaning con-

fused or obscured; or not enough to evaluate.

Language 30-27 Excellent to Very Good: Confident handling of appropriate constructions; hardly any errors of agreement, tense, number, word order, articles, pronouns, prepositions; meaning never obscured;

26-21 Good to Average: Well-formed sentences, but constructions not always appropriate to task (register); some errors of agreement, tense, number, word order, articles, pronouns, prepositions; meaning sometimes slightly obscured

20-15 Fair to Poor: Not sufficient variety of constructions available; problems mainly in complex constructions; several errors of negation, agreement, tense, number, word order/function, articles, pronouns, prepositions; meaning sometimes obscured.

14-9 Very Poor: Major problems in sentence construction; frequent errors of negation, agreement, tense, number, word order/function, articles, pronouns, prepositions; meaning often obscured.

Mechanics 5 Excellent to Very Good: Demonstrates mastery of conventions; hardly any errors

of spelling, punctuation, capitalization, paragraphing. 4 Good to Average: Occasional errors of spelling, punctuation, capitalization,

paragraphing 3-2 Fair to Poor: Frequent errors of spelling, punctuation, capitalization, paragraphing,

poor handwriting

Table 1: Sehr gut ≥ 90%, Gut ≥ 80%, Befriedigend ≥ 70%, Genügend ≥ 60%